answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What are the requirements of a probability distribution?

It is always non-negative.

The sum (or integral) over all possible outcomes is 1.

You roll 2 number cubes each labeled 1 to 6 what is probabilty you will roll a double?

The probability is based on the second cube having 6 sides. No matter what you roll on the first cube, the chance of rolling "doubles" is 1 in 6! Now if you ask, what is the chance in rolling a specific set of doubles, the chance is 1 in 6 times 1 in 6 times, or 1 in 36 ...

What is the probability of getting exactly 2 heads in 3 coin tosses?

With 3 coin tosses, the possible outcomes are:

HHH, HHT, HTH, HTT, THH, THT, TTH, TTT

There are 8 possible outcomes and 3 of them have 2 heads. Thus:

probability = 3/8 (= 0.375)

What is the probability of drawing a vowel from the word mathematics?

You have a total of 11 letters in "mathematics" and you have 4 vowels (a,e,a,i) so the probability of drawing a vowel is 4/11.

In other words if you were to consider the vowel's to be 1's and the consonants 2's. What is the probability of drawing a "1". There would be 4 1's and 7 2's. It would be 4/11

What is a biased spinner?

ok I'm sure not many people will be asking this but a biased spinner is basically a spinner that is unfair or not accurate

Why can't probability be negative?

Because probability is a measure of how likely it is that some event will occur. If there is no chance at all that the even will occur than the probability is 0. If it will occur with 100% certainty, then it is 1 or 100%

A negative number would have no meaning in terms of how likely an even is to occur?

You roll a standard number cube and get a multiple of 4?

You got a one or a four. One is the zeroeth multiple of 4. Four is the first multiple of four. Any other result, such as 8, and you are not rolling a standard die.

What are odds of probability of 9 over 11?

For 9 to 11 odds of winning;

Probability of winning:

0.55, or;

Chance of winning:

55%

The difference between an empirical distribution and a theoretical distribution?

Empirical Distribution: based on measurements that are actually taken on a variable.

Theoretical Distribution: not constructed on measurements but rather by making assumptions and representing these assumptions mathematically.

What is The difference between an empirical distribution and a theoretical distribution?

empirical distribution is based on your observation of out comes, it is based on real data. on the other hand theoretical is base on your theory regarding the distribution and the parameters, (i.e. normal/exponential...., u=5 vs u .5....and so on)

What is standard normal variate?

It's the same as a z-Transformation.

for all xi: (xi-mean(x)) / std(x)

Discrete and Continuous data is used in mathematics?

Yes it is. Discrete data is something that's set. Like say you were making a line graph about renting bikes. You can only rent whole bikes there is nothing in between. You shouldn't connect points on a line graph with discrete data but some cases can be argued.

Continuous data is usually a measurement that could change like time.

What is the probability of at least one birthday match among a group of 41 people?

The probability of at least 1 match is equivalent to 1 minus the probability of there being no matches.

The first person's birthday can fall on any day without a match, so the probability of no matches in a group of 1 is 365/365 = 1.

The second person's birthday must also fall on a free day, the probability of which is 364/365

The probability of the third person also falling on a free day is 363/365, which we must multiply by the probability of the second person's birthday being free as this must also happen. So for a group of 3 the probability of no clashes is (363*364)/(365*365).

Continuing this way, the probability of no matches in a group of 41 is (365*364*363*...326*325)/36541

This can also be written 365!/(324!*36541)

Which comes to 0.09685...

Therefore the probability of at least one match is 1 - 0.09685 = 0.9032

So the probability of at least one match is roughly 90%

What number will come up the most when rolling dice?

since there are 6 sides to it any of them could show up, 1,2,3,4,5, and 6 could all show up unlikely

How many different combinations of 4 books can be chosen from 6 books?


The number of 4 different book combinations you can choose from 6 books is;

6C4 =

6!/[4!(6-4)!] =

15 combinations of 4 different books.

What is a qualitative random variable?

A random variable which can take qualitative values rather than numeric values. For example, the question "What colour are your eyes?" will generate qualitative answers.

Trending Questions
What is the probability that you will spin the same number all 3 times but with 8 numbers? How many zeros are possible factorial 10 to the power factorial 10? Is In Clinical Trials Randomization Can Be Compared To rolling dice? What is the probability of P if you put all of the letters of probibility in a bag? Is the gaussian distribution discrete or continuous? What is the probability of rolling two dice and not getting the sum of two? What is represent outcome? Probability of winning the lottery? What is the probability of picking a picture card from a normal pack of 52 playing cards? What data cant be measured but divided into groups? If a box contains one red ball and one blue ball and two people touch one ball randomly with a one in two chance of picking either what is the chance of the blue ball being picked twice? What is the probability of flipping a coin three times and it landing on heads all three times? How does the genetic makeup of fraternal and identical twins differ? What is the probability of 12 number out of 37 on roulette not showing for 10 spins. 3725 time to the power of 10 50.42 What would be odds of this happening twice in a row with the same number? What value corresponds to the 38 percent of the data between the mean and the z value? Find Different arrangements that can be made out of the seven colours of rainbow so that blue and green colours shall not come together? If a coin is flipped 7 times what is the outcomes? Here is a spinner when the arrow is spun once a 1 2 or 3 can be scored bill is going to spin the arrow twice he will out his total score by adding the 2 scores he gets on the 2 spins work out x? Is there any event with probability one but no occurrence? What is the probability of rolling three dice and getting a two a four and a six?